What is the luminosity of a neutron star?

A neutron star has a radius of 10 km, and a temperature of 1 million K. What is its luminosity?
Answer #1

Answer: 0.2 solar luminosities


The formula to identify the luminosity of a star is,

L / L0 = (R/R0)2(T/T0)4

Here, R0 = 6.96*105 km

T0 = 5800 K

L / L0 = (10 km/6.96*105 km)2(1,000,000 K/5800 K)4

= 0.2 solar luminosities


answered by: anonymous